OpenStudy (anonymous):

what is the answer to this double integral? x= sqrt(y), x= 3 and y=0 , y= 9 sinPI*x^3 dxdy

OpenStudy (anonymous):

Change the order of integration to get \[\int\limits_{0}^{3} \int\limits_{0}^{x ^{2}}\sin( \pi x ^{3}) dy dx.\]

OpenStudy (anonymous):

Integrate the inside integral with respect to y first and evaluate between the limits 0 and x^2.

OpenStudy (anonymous):

oh. i knew it but i could fight new limits of integral. the fact is ive a test tomorrow and i dont know what this changing order of integral means. btw you solved it thank you. i know the rest

OpenStudy (anonymous):

could fight = couldnt find xd